OpenStudy (anonymous):

Which test point holds true for y − 2x ≤ 1? (0, 2) (-2, 4) (1, 4) (5, 0)

OpenStudy (anonymous):

substitute and check 2-0=2, not less than or equal to 1 4+8=12, not less than or equal to 1 4-2=2, nope 0-10=-10<1. the last one is correct

OpenStudy (anonymous):

thanks @GaryCollins